OpenStudy (anonymous):

Calculate the surface area of a cube with an edge length of 14cm

OpenStudy (anonymous):

6 faces Each face = 14 x 14cm 0= 196 cm^2 Therefore 6 x 196 cm^2 = 1176 cm^2

OpenStudy (anonymous):

1176 cm^2

OpenStudy (anonymous):

Sorry should read ... 6 faces Each face = 14 x 14cm = 196 cm^2 Therefore 6 x 196 cm^2 = 1176 cm^2

OpenStudy (anonymous):

A cube has six sides. 14cm x 14cm = 196cm^2 or 196 square centimeters. If each side has 196 square centimeters and we have 6 sides; then the total surface of the cube would be: 6 sides x 196 square centimeters or 1176 square centimeter

OpenStudy (anonymous):

area of one side=14^2=196 so total area=6*196=1176

OpenStudy (anonymous):

Shrishti I like your explanation but do we not refer to edges as sides? The question is looking atr the 6 faces of a cube. Just a thought )